OpenStudy (4meisu):

Solve the equation 3sin^2 x + 4 cos x - 4 = 0. x < 90 degrees

OpenStudy (anonymous):

Tip : write sin^2x = 1-cos^2x and then try to solve the quadratic

OpenStudy (lgbasallote):

did you get that @4meisu ?

OpenStudy (4meisu):

Kind of, not really though..

OpenStudy (anonymous):

3(1-cos^2x)+4cosx-4=0 3-3cos^2x+4cosx-4=0 let (cosx)=X then 3X^2+4X-1+0 then try to solve this and eleminat one of the tow solution ....

OpenStudy (4meisu):

Would the answer be cos =1/3 , 1?

OpenStudy (anonymous):

no that means that cosx=1/3 then search for x=cos^-1(1/3) search for ....you 'll get I dont remember just use a calculater to find it